Wildlife Co. reported net income of $8.3 million, interest expense of $0.5 million and $0.2 million of income tax expense. Wildlife's average total assets are $65.8 million and average stockholders' equity is $48.6 million. Wildlife's times interest earned ratio is closest to:

A. 6 times

B. 18 times

C. 41.5 times

D.45 times
Answer #1

Interest earned ratio = Income before interest and taxes / Interest expense

Income before interest and taxes = 8.3 + 0.5 + 0.2 = 9

Interest earned ratio = 9 / 0.5 = 18 times

Option B is correct.